### v1

@@ 1 Introduction
Functional data is commonly observed in a wide range of fields, and there has been significant research on functional data analysis (FDA) with applications in diverse areas such as biomedical studies, neuroscience, linguistics, psychology, demography, economics, and engineering. For a comprehensive review of the FDA and its applications, we recommend referring to Ramsay and Silverman, [29] and Wang et al., [37].
Estimating the mean function is a fundamental problem in the FDA that has garnered considerable attention in the existing literature. Notable references include Rice and Silverman, [31], Page et al., [25], Park et al., [27], Jiang et al., [16], and Cai and Yuan, [9]. This problem arises naturally in a broad range of practical applications, such as analyzing a diffusion tensor imaging (DTI) dataset to investigate multiple sclerosis (MS) patients [33, 12, 28], analyzing a longitudinal CD4 cell count dataset to investigate acquired immunodeficiency syndrome (AIDS) [42], and examining a longitudinal dataset of trajectories of patient-reported symptom severity of Parkinson’s disease [38].
On the other hand, transfer learning has experienced growing popularity as a machine learning technique aimed at improving performance in a target domain by utilizing information from different yet related source domains. This approach proves particularly valuable in situations where acquiring target observations is infrequent or costly, but data from analogous studies are available. Transfer learning has been applied in various machine learning applications, including computer vision [35, 13], speech recognition [14], and genre classification [11]. We refer readers to Pan and Yang, [26] and Weiss et al., [39] for a more comprehensive discussion on the practical applications of transfer learning. Recently, transfer learning has also attracted increasing attention in statistics and has been successful in various statistical learning problems such as nonparametric classification [7, 30], high-dimensional linear regression [21], large-scale Gaussian graphical model [22], high-dimensional generalized linear models [23, 34], nonparametric regression [6], and contextual multi-armed bandits [3].
@@
An interesting and motivating example of functional mean estimation with transfer learning is found in the trend analysis of coronavirus disease-19 (COVID-19). The objective is to gain insights into the epidemic’s progression and identify overall trends in the United States. Similar to the work by Kalogridis and Van Aelst, [17], one can look at the number of new daily COVID-19 cases across the 50 U.S. states. Each state’s curve may be regarded as a functional realization and the problem reduces to estimate their functional mean. However, COVID-19 has had a global impact, affecting populations worldwide [18]. Structural similarities in trends across different countries have been observed, as demonstrated by James and Menzies, [15]. This suggests that transfer learning can improve the estimation accuracy of the COVID-19 trend in the United States by utilizing the COVID-19 curves from other American countries or other continents, including Africa, Asia, Europe, and Oceania.
@@ 1.1 Problem formulation
where $T^{[t]}_{ij}$ are target design points, and $\varepsilon^{[t]}_{ij}$ are an independent random noises. Based on the target sample $\mathcal{D}^{[t]}:=\{(T_{ij}^{[t]},Y_{ij}^{[t]}):j=1,\ldots,m_{i}^{[t]},~{}i=1,\ldots,n^{[t]}\}$, our primary objective is to estimate the target mean function $f^{[t]}(\cdot):=\mathbb{E}(X^{[t]}(\cdot))$. Such problems naturally arise in a wide range of applications and are typical in FDA. Extensive examples can be found in the works of Ramsay and Silverman, [29].
@@ 1.2 Our contribution
We turn our attention to the low sampling frequency regime where $m_{t}\ll n_{t}^{1/2\alpha_{m}}$ holds. In this regime, we shall explore distinct conditions under which transfer learning can enhance performance under common and independent designs, respectively. Under a common design, transfer learning proves effective when $m_{s}^{-{2\alpha_{m}}}+(Kn_{s})^{-1}\ll m_{t}^{-{2\alpha_{m}}}$, whereas it is effective if $(Km_{s}n_{s})^{-2\alpha_{m}/(2\alpha_{m}+1)}+(Kn_{s})^{-1}\ll(m_{t}n_{t})^{-2\alpha_{m}/(2\alpha_{m}+1)}$ under an independent design. To a certain extent, both designs require a sufficient number of total subjects in the source samples $(Kn_{s}\gg m_{t}^{2\alpha_{m}}~{}\text{and}~{}Kn_{s}\gg(m_{t}n_{t})^{2\alpha_{m}/(2\alpha_{m}+1)},~{}\text{respectively})$, but neither implies the other. The second condition, however, exhibits a significant disparity between the two designs. In a common design, a higher sampling frequency in the source samples ($m_{s}\gg m_{t}$) is necessary, while in an independent design, a larger total number of observations in the source samples ($Km_{s}n_{s}\gg m_{t}n_{t}$) is required. Finally, another noteworthy observation is that the size of the source group $K$ plays a more influential role in an independent design than in a common design. This mainly is because increasing $K$ leads to a larger total number of observations in the source samples, but not to a higher sampling frequency in the source samples.
@@
Although our primary focus is on transfer learning, these results also provide new insights into functional mean estimation in the conventional setup. By setting both the source sample size $n_{s}$ and sampling frequency $m_{s}$ to zero, we obtain rates for functional mean estimation without transfer learning under both common and independent designs. Our rates of convergence, $m_{t}^{-2\alpha_{m}}+n_{t}^{-1}$ and $(m_{t}n_{t})^{-2\alpha_{m}/(2\alpha_{m}+1)}+n_{t}^{-1}$, serve as a generalization of the rates derived by Cai and Yuan, [9]. This generalization is particularly significant when the target mean function is rough and even not differentiable, as the current framework still guarantees consistent estimation. Such flexibility is a key advantage of our model.
@@ 2 Transfer learning for functional mean estimation under a common design
The random functions and noises are uniformly sub-Gaussian variables with a positive variance proxy $\tau^{2}$. As per Wainwright, [36], we assume that for every $u>0$, the following holds: for any given $x\in[0,1]$ and $k=1,\ldots,K$,
@@ 2.1 Methodology and upper bound
In this subsection, we present a novel algorithm for estimating the target mean function. Before delving into the transfer learning problem of our main interest, it is useful to revisit the conventional framework for functional mean estimation, which is a crucial component of the transfer learning algorithm. This involves estimating the target mean function $f^{[t]}$ when no source samples are available, i.e. $n_{s}=0$. Although this problem has been extensively investigated by Cai and Yuan, [9], and the focus of this paper is on transfer learning, it is still valuable to discuss it separately.
@@
Theorem 2.1 generalizes the minimax rate shown by Cai and Yuan, [9], $\Theta(m_{t}^{-2r}+n_{t}^{-1})$, where both the curve $X^{[t]}$ and mean function $f^{[t]}$ are assumed to be $r$ times differentiable with $r\in\mathbb{Z}^{+}$. Notably, the current framework remains consistent in estimating the mean function even when it is not smooth $(\alpha_{m}<1)$, whereas the previous one does not. Moreover, our approach provides greater flexibility by not requiring the smoothness of the target curve $X^{[t]}$. Therefore, the proposed model represents a significant improvement over the existing literature on functional mean estimation.
@@
3:Let us denote by $\mathcal{D}^{(i)}:=\{(T_{ij},Y_{ij}):j=1,\ldots,m\}$ the collection of observations from the same subject $i\in\{1,\ldots,n\}$. For index $r=1,\ldots,b^{-1}$, we take randomized collection $\mathcal{D}_{r}:=\{({\mathbb{T}}_{i,r},\mathbb{Y}_{i,r}):i=1,\ldots,n\}$ of observations where $(\mathbb{T}_{i,r},\mathbb{Y}_{i,r})$ is randomly chosen from the following process:
@@
Consider any $(1/m)$-packing and $(1/2m)$-covering sub-collection $\widetilde{\mathcal{D}}^{(i)}$ of $\mathcal{D}^{(i)}$ where the distance between observations is computed based on design points. The random $(\mathbb{T}_{i,r},\mathbb{Y}_{i,r})$ is now uniformly chosen from $\{(T,Y)\in\widetilde{\mathcal{D}}^{(i)}:T\in I_{r}\}$.
@@
7:     The random $(\mathbb{T}_{i,r},\mathbb{Y}_{i,r})$ is uniformly chosen from $\{(T,Y)\in\mathcal{D}^{(i)}:T\in I_{r}\}$.
@@
[…]
@@
[…]
@@ 2.3 Adaptive estimation
Notice that the base algorithms $\mathcal{A}_{\mathrm{TL}}$ and $\mathcal{A}_{\mathrm{CL}}$ generates a randomized estimator. Consequently, the adaptive learning algorithm $\mathcal{A}_{\mathrm{ALC}}$ must be randomized as well. To improve the performance in finite samples, we execute this adaptive procedure $r_{\mathrm{max}}$ times and calculate the average of the resulting estimates. This approach shares a similar idea with bagging estimators suggested by Breiman, [2]. The optimal minimax rate of convergence, up to logarithmic terms, is achieved by this algorithm, as demonstrated by the subsequent theorem.
@@ 3.1 Methodology and upper bound
The derived minimax rate of convergence again generalizes the optimal rate shown by Cai and Yuan, [9], $\Theta((m_{t}n_{t})^{-2r/(2r+1)}+n_{t}^{-1})$, where they assumed that both the curve $X^{[t]}$ and the mean function $f^{[t]}$ are differentiable $r$ times $(r\in\mathbb{Z}^{+})$. Our framework provides greater flexibility and improvement over previous approaches for functional mean estimation, similar to what we have observed under a common design.
@@
[…]
@@ 3.3 Adaptive estimation
Similar to the adaptive learning algorithm $\mathcal{A}_{\mathrm{ALC}}$ for a common design, the adaptive algorithm $\mathcal{A}_{\mathrm{ALI}}$ for an independent design also involves multiple steps of random down-sampling, resulting in a randomized estimator. To improve the performance of estimation in finite samples, we will execute this adaptive process $r_{\mathrm{max}}$ times and average the resulting estimates. This approach is analogous to bagging estimators proposed by Breiman, [2]. As demonstrated by the subsequent theorem, this algorithm achieves the optimal minimax rate of convergence, up to logarithmic factors.
@@ 5 Discussion
Another important problem is transfer learning for functional linear regression. In the conventional settings, Yuan and Cai, [41] established the minimax rate of convergence for estimating the regression function, while both Cai and Hall, [4] and Cai and Yuan, [10] investigated the optimal prediction problem. It is of significant interest to explore transfer learning for linear regression and investigate the conditions under which source samples can improve the estimation performance.
@@
Finally, in the independent design setting, we assumed that the marginal distribution of the source samples only trivially differs from that of the target sample. In practice, the marginal density ratio between the target and source samples may not be uniformly bounded away from zero and infinity. This departure from the uniformity of density ratio can significantly impact the minimax rate of convergence. For example, if the marginal distributions of the target and source samples are completely singular, the benefit of transfer learning may be negligible. Incorporating this effect into the analysis could be achieved by introducing a new measure to quantify the singularity between marginal distributions. Similar to the transfer exponent introduced in Kpotufe and Martinet, [19], such a measure would be a valuable addition to assess the influence of marginal singularity under the transfer learning setup.
@@ 6 Proofs
We shall prove the main theorems regarding the common design in this section, including conventional learning (Theorem 2.1), transfer learning (Theorem 2.2 and 2.3) and adaptive transfer learning (Theorem 2.4). The proofs of main theorems associated with the independent design and every other technical result are presented in the Supplementary Material [5].
Before diving into the proof, let us briefly introduce some notations to be used in the proof. $\lVert x\rVert_{2}$ denotes the $\ell_{2}$-norm and $\lVert x\rVert_{\infty}$ denotes the $\ell_{\infty}$-norm for any vector $x\in\mathbb{R}^{k}$. In addition, for any square matrix $A\in\mathbb{R}^{k\times k}$, we denote the Frobenius norm by $\lVert A\rVert_{F}$ and $\ell_{2}$-norm by $\lVert A\rVert_{2}$. On the other hand, for any probability measures $\mathbb{P}$ and $\mathbb{Q}$, the total variation distance is denoted as $\lVert\mathbb{P}-\mathbb{Q}\rVert_{\mathrm{TV}}$ and the Kullback-Leibler divergence as $D_{\mathrm{KL}}(\mathbb{P}\,\|\,\mathbb{Q})$.

### v2

@@ Abstract
62J05,
62G20,
Adaptivity,
common design,
functional data analysis,
independent design,
mean function,
minimax rate of convergence,
phase transition,
transfer learning,
keywords:
[class=MSC]
keywords:
[…]
, , and
[…]
@@ 1 Introduction
Functional data is commonly observed in a wide range of fields, and there has been significant research on functional data analysis (FDA) with applications in diverse areas such as biomedical studies, neuroscience, linguistics, psychology, demography, economics, and engineering. For a comprehensive review of the FDA and its applications, we recommend referring to Ramsay and Silverman [29] and Wang, Chiou and Müller [37].
Estimating the mean function is a fundamental problem in the FDA that has garnered considerable attention in the existing literature. Notable references include Rice and Silverman [31], Page et al. [25], Park et al. [27], Jiang, Aston and Wang [16], and Cai and Yuan [9]. This problem arises naturally in a broad range of practical applications, such as analyzing a diffusion tensor imaging (DTI) dataset to investigate multiple sclerosis (MS) patients [33, 12, 28], analyzing a longitudinal CD4 cell count dataset to investigate acquired immunodeficiency syndrome (AIDS) [42], and examining a longitudinal dataset of trajectories of patient-reported symptom severity of Parkinson’s disease [38].
On the other hand, transfer learning has experienced growing popularity as a machine learning technique aimed at improving performance in a target domain by utilizing information from different yet related source domains. This approach proves particularly valuable in situations where acquiring target observations is infrequent or costly, but data from analogous studies are available. Transfer learning has been applied in various machine learning applications, including computer vision [35, 13], speech recognition [14], and genre classification [11]. We refer readers to Pan and Yang [26] and Weiss, Khoshgoftaar and Wang [39] for a more comprehensive discussion on the practical applications of transfer learning. Recently, transfer learning has also attracted increasing attention in statistics and has been successful in various statistical learning problems such as nonparametric classification [7, 30], high-dimensional linear regression [21], large-scale Gaussian graphical model [22], high-dimensional generalized linear models [23, 34], nonparametric regression [6], and contextual multi-armed bandits [3].
@@
An interesting and motivating example of functional mean estimation with transfer learning is found in the trend analysis of coronavirus disease-19 (COVID-19). The objective is to gain insights into the epidemic’s progression and identify overall trends in the United States. Similar to the work by Kalogridis and Van Aelst [17], one can look at the number of new daily COVID-19 cases across the 50 U.S. states. Each state’s curve may be regarded as a functional realization and the problem reduces to estimate their functional mean. However, COVID-19 has had a global impact, affecting populations worldwide [18]. Structural similarities in trends across different countries have been observed, as demonstrated by James and Menzies [15]. This suggests that transfer learning can improve the estimation accuracy of the COVID-19 trend in the United States by utilizing the COVID-19 curves from other American countries or other continents, including Africa, Asia, Europe, and Oceania.
@@ 1.1 Problem formulation
where $T^{[t]}_{ij}$ are target design points, and $\varepsilon^{[t]}_{ij}$ are an independent random noises. Based on the target sample $\mathcal{D}^{[t]}:=\{(T_{ij}^{[t]},Y_{ij}^{[t]}):j=1,\ldots,m_{i}^{[t]},~{}i=1,\ldots,n^{[t]}\}$, our primary objective is to estimate the target mean function $f^{[t]}(\cdot):=\mathbb{E}(X^{[t]}(\cdot))$. Such problems naturally arise in a wide range of applications and are typical in FDA. Extensive examples can be found in the works of Ramsay and Silverman [29].
@@ 1.2 Our contribution
We turn our attention to the low sampling frequency regime where $m_{t}\ll n_{t}^{1/2\alpha_{m}}$ holds. In this regime, we shall explore distinct conditions under which transfer learning can enhance performance under common and independent designs, respectively. Under a common design, transfer learning proves effective when $m_{s}^{-{2\alpha_{m}}}+(Kn_{s})^{-1}\ll m_{t}^{-{2\alpha_{m}}}$, whereas it is effective if $(Km_{s}n_{s})^{-2\alpha_{m}/(2\alpha_{m}+1)}+(Kn_{s})^{-1}\ll(m_{t}n_{t})^{-2\alpha_{m}/(2\alpha_{m}+1)}$ under an independent design. To a certain extent, both designs require a sufficient number of total subjects in the source samples $(Kn_{s}\gg m_{t}^{2\alpha_{m}}~{}\text{and}~{}Kn_{s}\gg(m_{t}n_{t})^{2\alpha_{m}/(2\alpha_{m}+1)},~{}\text{respectively})$, but neither implies the other. The second condition, however, exhibits a significant disparity between the two designs. In a common design, a higher sampling frequency in the source samples ($m_{s}\gg m_{t}$) is necessary, while in an independent design, a larger total number of observations in the source samples ($Km_{s}n_{s}\gg m_{t}n_{t}$) is required. Finally, another noteworthy observation is that the size of the source group $K$ plays a more influential role in an independent design than in a common design. This is mainly because increasing $K$ leads to a larger total number of observations in the source samples, but not to a higher sampling frequency in the source samples.
@@
Although our primary focus is on transfer learning, these results also provide new insights into functional mean estimation in the conventional setup. By setting both the source sample size $n_{s}$ and sampling frequency $m_{s}$ to zero, we obtain rates for functional mean estimation without transfer learning under both common and independent designs. Our rates of convergence, $m_{t}^{-2\alpha_{m}}+n_{t}^{-1}$ and $(m_{t}n_{t})^{-2\alpha_{m}/(2\alpha_{m}+1)}+n_{t}^{-1}$, serve as a generalization of the rates derived by Cai and Yuan [9]. This generalization is particularly significant when the target mean function is rough and even not differentiable, as the current framework still guarantees consistent estimation. Such flexibility is a key advantage of our model.
@@ 2 Transfer learning for functional mean estimation under a common design
The random functions and noises are uniformly sub-Gaussian variables with a positive variance proxy $\tau^{2}$. As per Wainwright [36], we assume that for every $u>0$, the following holds: for any given $x\in[0,1]$ and $k=1,\ldots,K$,
@@ 2.1 Methodology and upper bound
In this subsection, we present a novel algorithm for estimating the target mean function. Before delving into the transfer learning problem of our main interest, it is useful to revisit the conventional framework for functional mean estimation, which is a crucial component of the transfer learning algorithm. This involves estimating the target mean function $f^{[t]}$ when no source samples are available, i.e. $n_{s}=0$. Although this problem has been extensively investigated by Cai and Yuan [9], and the focus of this paper is on transfer learning, it is still valuable to discuss it separately.
@@
Theorem 2.1 generalizes the minimax rate shown by Cai and Yuan [9], $\Theta(m_{t}^{-2r}+n_{t}^{-1})$, where both the curve $X^{[t]}$ and mean function $f^{[t]}$ are assumed to be $r$ times differentiable with $r\in\mathbb{Z}^{+}$. Notably, the current framework remains consistent in estimating the mean function even when it is not smooth $(\alpha_{m}<1)$, whereas the previous one does not. Moreover, our approach provides greater flexibility by not requiring the smoothness of the target curve $X^{[t]}$. Therefore, the proposed model represents a significant improvement over the existing literature on functional mean estimation.
@@
[…]
@@
[…]
@@
[…]
@@
$\max_{1\leq j\leq m_{t}+1}\bigl{(}T_{j}^{[t]}-T_{j-1}^{[t]}\bigr{)}\leq\frac{C_{t}}{m_{t}}\qquad\text{and}\qquad\max_{1\leq j\leq m_{s}+1}\bigl{(}T_{j}^{[s]}-T_{j-1}^{[s]}\bigr{)}\leq\frac{C_{s}}{m_{s}}$
@@
$\widehat{f}^{[t]}=\begin{dcases}\widehat{f}_{\mathrm{CL}}^{[t]}&\text{when}~{}R_{\mathrm{C}}(\widehat{f}_{\mathrm{CL}}^{[t]})\leq R_{\mathrm{C}}(\widehat{f}_{\mathrm{TL}}^{[t]}),\\ \widehat{f}_{\mathrm{TL}}^{[t]}&\text{when}~{}R_{\mathrm{C}}(\widehat{f}_{\mathrm{CL}}^{[t]})>R_{\mathrm{C}}(\widehat{f}_{\mathrm{TL}}^{[t]}),\end{dcases}$
@@ 2.3 Adaptive estimation
Notice that the base algorithms $\mathcal{A}_{\mathrm{TL}}$ and $\mathcal{A}_{\mathrm{CL}}$ generates a randomized estimator. Consequently, the adaptive learning algorithm $\mathcal{A}_{\mathrm{ALC}}$ must be randomized as well. To improve the performance in finite samples, we execute this adaptive procedure $r_{\mathrm{max}}$ times and calculate the average of the resulting estimates. This approach shares a similar idea with bagging estimators suggested by Breiman [2]. The optimal minimax rate of convergence, up to logarithmic terms, is achieved by this algorithm, as demonstrated by the subsequent theorem.
@@ 3.1 Methodology and upper bound
The derived minimax rate of convergence again generalizes the optimal rate shown by Cai and Yuan [9], $\Theta((m_{t}n_{t})^{-2r/(2r+1)}+n_{t}^{-1})$, where they assumed that both the curve $X^{[t]}$ and the mean function $f^{[t]}$ are differentiable $r$ times $(r\in\mathbb{Z}^{+})$. Our framework provides greater flexibility and improvement over previous approaches for functional mean estimation, similar to what we have observed under a common design.
@@
$\widehat{f}^{[t]}=\begin{dcases}\widehat{f}_{\mathrm{CL}}^{[t]}&\text{when}~{}R_{\mathrm{I}}(\widehat{f}_{\mathrm{CL}}^{[t]})\leq R_{\mathrm{I}}(\widehat{f}_{\mathrm{TL}}^{[t]}),\\ \widehat{f}_{\mathrm{TL}}^{[t]}&\text{when}~{}R_{\mathrm{I}}(\widehat{f}_{\mathrm{CL}}^{[t]})>R_{\mathrm{I}}(\widehat{f}_{\mathrm{TL}}^{[t]}),\end{dcases}$
@@ 3.3 Adaptive estimation
Similar to the adaptive learning algorithm $\mathcal{A}_{\mathrm{ALC}}$ for a common design, the adaptive algorithm $\mathcal{A}_{\mathrm{ALI}}$ for an independent design also involves multiple steps of random down-sampling, resulting in a randomized estimator. To improve the performance of estimation in finite samples, we will execute this adaptive process $r_{\mathrm{max}}$ times and average the resulting estimates. This approach is analogous to bagging estimators proposed by Breiman [2]. As demonstrated by the subsequent theorem, this algorithm achieves the optimal minimax rate of convergence, up to logarithmic factors.
@@ 5 Discussion
Another important problem is transfer learning for functional linear regression. In the conventional settings, Yuan and Cai [41] established the minimax rate of convergence for estimating the regression function, while both Cai and Hall [4] and Cai and Yuan [10] investigated the optimal prediction problem. It is of significant interest to explore transfer learning for linear regression and investigate the conditions under which source samples can improve the estimation performance.
@@
Finally, in the independent design setting, we assumed that the marginal distribution of the source samples only trivially differs from that of the target sample. In practice, the marginal density ratio between the target and source samples may not be uniformly bounded away from zero and infinity. This departure from the uniformity of density ratio can significantly impact the minimax rate of convergence. For example, if the marginal distributions of the target and source samples are completely singular, the benefit of transfer learning may be negligible. Incorporating this effect into the analysis could be achieved by introducing a new measure to quantify the singularity between marginal distributions. Similar to the transfer exponent introduced in Kpotufe and Martinet [19], such a measure would be a valuable addition to assess the influence of marginal singularity under the transfer learning setup.
@@ 6 Proofs
This section contains the proofs of our main results for common design, especially under transfer learning setup: Theorem 2.2 and Theorem 2.3. We will defer the proofs of the remaining results for common design, as well as the results for independent design and every other technical results, to the Supplementary Material [5].